An experienced carpenter can panel a room 3 times faster than an apprentice can. Working together, they can panel the room in 6 hours. How long would it take each person working alone to do the job?: An experienced carpenter can panel a room 3 times faster than an apprentice can. Working together, they can panel the room in 6 hours. How long would it take each person working alone to do the job?
Answer by ptaylor(1326)


Let x=the rate at which the apprentice can panel the room
Then 3x=rate at which the carpenter can panel the room
Together the two can panel the room at the rate of x+3x or 4x rooms per hour.
So, our equation to solve is:
4x rooms per hour times 6 hours equals one room or:
4x*6=1
24x=1
x=1/24 rooms per hour -----rate of the apprentice
So, in 24 hours, the apprentice can panel the room
3x=3/24 rooms per hour=1/8 rooms per hour-----rate of the carpenter
So in 8 hours the carpenter can panel the room
together they panel at the rate of 1/24 + 3/24 or 4/24=1/6 rooms per hour which means that in 6 hours they can panel one room
Hope this helps-----ptaylor